OpenStudy (knightofpentacles):

600 at 5% for 2 years how to do the problem step by step please demonstrate and thank you very immensely appreciated.

OpenStudy (knightofpentacles):

process matters most xD

OpenStudy (dan815):

xD what does it even mean 600 at 5%

OpenStudy (dan815):

like you deposited 600, and theres a return of 5% a year?

OpenStudy (dan815):

or compounded 5%

OpenStudy (knightofpentacles):

simple interest

OpenStudy (dan815):

okay so u will get 5% of 600 every year

OpenStudy (knightofpentacles):

principal annual interest rate and time

OpenStudy (knightofpentacles):

ok

OpenStudy (dan815):

you will get 5% , twice as there are 2 years so 2*(5% of 600)=2*0.05 * 600 = total interest in 2 years

OpenStudy (knightofpentacles):